OpenStudy (anonymous):

Pressing the brake of a car causes it to slow down from 38 m/s at a rate of -2.0 m/s2 for a period of 4 s. How far did the car travel during these 4 s?

OpenStudy (anonymous):

Remember the kinematic equations ^_^ \[s=v_0t + at^2\] \[v_0=38m/s \\ a=-2m/s^2 \\ t=4s\]

OpenStudy (anonymous):

isn't this the equation for speed?

OpenStudy (anonymous):

First of all, there's a typo in the equation. It should be (1/2)at^2. Sorry about that. Second, "s" is displacement in this case - it can also be written as "x" (if it's moving in the x direction) to avoid confusion \[x=v_0t + \frac{1}{2}at^2\]

OpenStudy (anonymous):

You can tell by looking at the units. we want units of distance for the answer, and on the right hand side, we have velocity multiplied by time \[\text{velocity} \times \text{time} =\left( \frac{\text{distance}}{\text{time}} \right) \times \text{time} = \text{distance}\] and acceleration multiplied by time squared \[ \text{acceleration} \times (\text{time})^2 = \left( \frac{\text{distance}}{(\text{time})^2} \right) \times (\text{time})^2 = \text{distance}\]